OpenStudy (anonymous):

E (0,5), F ( -2,-1), and G ( 4,1) are vertices of a triangle. a) find the coordinates of the midpoint of EF (call it J) and EG (call it K) b) show that JK is parallel to FG c) show that JK is half the length of FG

OpenStudy (mertsj):

Do you know how to find the midpoint?

OpenStudy (anonymous):

yah can yu tell me if the ans is right though i got (-1,2) for the midpoint of EF

OpenStudy (mertsj):

that is correct! Good job!

OpenStudy (anonymous):

and the midpoint of EG is (2,3)

OpenStudy (anonymous):

im not sure wht to do for b and c

OpenStudy (mertsj):

To show that the two are parallel, show that they have the same slopes.

OpenStudy (anonymous):

uhm i got 1/3

OpenStudy (anonymous):

:/

OpenStudy (mertsj):

For each one?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i got the same slop now how abt c)

OpenStudy (mertsj):

To show that c is true, find the length of each segment and show that one is twice the other. Use the distance formula.

OpenStudy (anonymous):

so wait.. the length of JK and FG?

OpenStudy (mertsj):

Yes. Because the problem says to show that JK is half of FG.

OpenStudy (anonymous):

i got 10 as the length for jk can yu try it and tell me if its right

OpenStudy (mertsj):

Hang on.

OpenStudy (mertsj):

I believe you forgot to take the square root. I got sqrt10

OpenStudy (anonymous):

oh which is 3.16

OpenStudy (mertsj):

I would just leave it sqrt10 and then find FG. I bet it will turn out to be 2sqrt10

OpenStudy (anonymous):

i actually got 40 :/

OpenStudy (mertsj):

Did you remember to take the square root?

OpenStudy (anonymous):

6.32

OpenStudy (anonymous):

oh thanks :P

OpenStudy (mertsj):

\[\sqrt{40}=\sqrt{4(10)}=2\sqrt{10}\]

OpenStudy (mertsj):

yw